Please help me with this problem:
x^2+8x=12
I think the finished answer has to be written with a sqrt or 2 answers.

Answer by stanbon(75887)

x^2+8x=12
-----
x^2 + 8x - 12 = 0
---
x = [-8 +- sqrt(64-4*-12)]/2
----
x = [-8 +- sqrt(112)]/2
---
x = [-8 +- 4sqrt(7)]/2
------
x = -4+2sqrt(7) or x = -4-2sqrt(7)
==========================
Cheers,
Stan H.
==================